# A reduced-order framework for temperature estimation in food freezing from optimally located sensors, including turbulent conjugate flow scenarios

### v1

@@ 4.3 Greedy algorithm for optimal sensor placement
$\frac{\|T-T^{*}\|}{T}\leq\left(\inf_{c\in\mathbb{R}^{n}}\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{-1}\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}}$ (12)
where $\boldsymbol{A}_{n}$ is the $n$-th low rank approximation of $\boldsymbol{A}$ and ${\|\cdot\|_{\text{FRO}}}$ denotes the Frobenius norm [61]:
@@
The bound (12) shows where the condition $m>n$ comes from, indicating that a reduced model that is too enriched would lead to non-observable modes, a critical issue for well-posedness in data assimilation problems. Using standard arguments, the second term in the right-hand-side of (12) can be expressed in terms of the singular values of $\boldsymbol{A}$:
$\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}}=\frac{\sqrt{\sum_{i=n+1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}{\sqrt{\sum_{i=1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}$
@@
$\left(\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{2}$ $=\frac{d^{T}\hat{V}^{T}\boldsymbol{G}^{T}\boldsymbol{G}\hat{V}d}{d^{T}V^{T}\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}Vd}=\frac{d^{T}\hat{S}^{2}d}{d^{T}d}=\frac{\sum_{i=1}^{n}d_{i}d_{j}\hat{S}_{i}^{2}}{\sum_{i=1}^{n}d_{i}^{2}}$ (13)
@@
Then, for a given $n$, we aim at selecting the location of $m\geq n$ sensors from a pool of $M$ possible choices over the observable region $\Omega$. The proposed approach follows the underlying idea of a Greedy method, adding the sensors progressively and looking at each stage for the location of a new sensor that yields the smallest a-priori bound (14). Let us recall that, as seen in 4, a measurement sensor can be seen as a linear operation on the temperature field. Thus, considering the Riesz representation of this linear functional, in what follows, we will refer to a sensor as a column of the matrix $\boldsymbol{W}$. Moreover, we will denote as $\mathcal{S}=\{\ell_{1},\ldots,\ell_{M}\}$ a set of all possible sensor locations. The step-by-step process is shown in the blue box (Algorithm 1). Besides optimizing the conditioning of the matrix in the normal equations, this approach for the selection of additional modes has the physical meaning of adding experimental measurements at locations where the most relevant dynamics can be observed.
@@ 5 Numerical examples
The pipeline of the reduced-order inverse estimation framework proposed in this work is illustrated in Figure 8. In the training stage, we construct a five-dimensional parameter set by gathering snapshots from 64 simulations of the direct model. The resulting time-dependent dataset is obtained by randomly varying the boundary conditions, i.e., $U_{in}$, $T_{C}$, $T_{ext}$, and $h_{ext}$, within a fixed range by using a uniform random distribution, as presented in the lower-left corner of Figure 8. The magnitude of these variation ranges is consistent with real operating conditions of convective freezers and was estimated from the literature review and technical information from previous works [49, 23]. Then, we train the reduced order space by using an SVD with 48 simulations from the previous dataset, resulting in a total of 1.2×$\times\dot{1}0^{4}$ snapshots. The first 15 modes of the resulting bases are depicted in Figure 9. Thereafter, we manufacture artificial low-quality measurements of the temperature fields from the direct model by degrading the resolution in sensors or pixels of size $2\times 2$ cm2 (see also Figure 7). Finally, using a Greedy algorithm, we find optimal sensor placements based on a-priori error optimization. In the reconstruction stage, we recover the time-dependent temperature fields from 16 synthetic measurements using direct simulations different from those within the training dataset, using the ROM-based least-squares approach described in the previous section.
For the computational implementation of the numerical experiments, for the finite volume scheme solution of the forward problem, we have used an in-house library called MVF2D [62]. For the full ROM and inverse pipeline, we have used the software MAD ([63, 64], chapter 5), used successfully for forward and inverse problems in previous works such as [42, 65, 66].
Figure 8: Flow diagram of the present inverse framework used to reconstruct the temperatures.
Figure 9: The first 15th ordered (from left to right and top to bottom) modes of the reduced-order model for the direct problem using 48 simulations and a singular value decomposition.
@@ 5.1 Reconstruction of temperature fields from thermocamera measurements
The first test considers full-domain synthetic measurements mimicking a thermocamera placed over the two-dimensional mid-plane of the freezer. This setup results in an image composed of $m=361$ pixels or sensors. Having this setup fixed for the measurement vectors $\ell_{i}$, one can evaluate the matrix conditioning of the normal equations (10) summarized in (14). The decay of the eigenvalues for the covariance matrix associated with the snapshot matrix is shown in Figure 10. The values and decay tendency of the eigenvalues are related to the ROM’s ability to approximate the original model’s physics. The rapid decay of the values suggests, in particular, that the SVD provides a good approximation of the temperature field in the phase change phenomenon, also with a limited number of modes.
The indicator $e(n)$ from (14) for the fixed measurement setup is shown in Figure 10b. Since this quantity is related to an upper bound for the average error when computing the reconstruction, we infer an optimal dimension for the reduced-order model of $n=111$. The curve also clearly shows how adding more modes affects the matrix conditioning, lowering the quality of the expected result due to the reduced observability of the ROM when $n$ approaches $m$.
(a) Eigen-values decay for $\boldsymbol{A}^{T}\boldsymbol{A}$
(b) A-priori error (14)
Figure 10: Reconstruction error study for 2D food-freezing numerical example. Optimal results are achieved according to this bound for $n=111$.
A comparison of the temperature fields at three different snapshots is provided in Figure 12, showing the ground truth $T_{\text{GT}}$ (not included in the reduced-order model training set), the measurements $\ell=\boldsymbol{W}^{T}T_{\text{GT}}$, and the reconstructions $T^{*}$. As a benchmark criteria, we introduce the relative $\ell^{2}$ error:
@@
which we evaluate for every test case. The average result for the 16 simulations is depicted with respect to time in Figure 11, where we can conclude an overall good reconstruction quality during the whole freezing process, thus validating our framework at a proof-of-concept stage.
In addition, we do test exhaustively the reconstruction algorithm using a single, randomly chosen, parameter configuration, with the following values: $U_{in}=0.2\ ms^{-1}$, $T_{C}=-21.7\ \degree C$, $T_{ext}=21.3\ \degree C$, and $h_{ext}=0.88\ Wm^{-2}K^{-1}$.
Figure 11: Temporal evolution of the average reconstruction error for the full test dataset of 16 simulations out of the training set.
We observe a good qualitative agreement between the ground truth and the reconstructed field during the whole freezing process, especially within the food domain, which is of utmost interest for the targeted application. The field reconstruction indicates that the turbulent airflow recirculation through the cabinet is properly described. Moreover, the congruence in the concentration of freezing fronts within the food demonstrates that the phase change process is adequately represented. However, slight differences are noticeable in the upper region of the cabinet, where the airflow is more chaotic due to turbulent fluctuations and the stratification by buoyant forces.
Figure 12: Temperature field evolution of (i) the ground-truth from the direct model, (ii) the full-domain measurements, and (iii) the reconstruction procedure at three characteristic instants of time.
@@
$\Phi_{rel,error}(\%)=\frac{|\Phi_{P}-\Phi^{*}_{P}|}{|\overline{\Phi}_{P}|}\times 100$ (16)
@@
By analyzing the time-dependent relative errors of temperature and freezing rate (see Fig. 13, bottom), we find that the control point in the freezing center presents larger errors than in the food boundaries for the whole process. We must consider that the freezing center represents the point where the freezing fronts concentrate after passing through the entire food domain, so it accumulates all the numerical errors in space and time of the nonlinear food phase-change phenomenon. Therefore, it can be considered that this area is prone to greater numerical errors, which could explain the lower quality reconstruction. Moreover, the local errors increase mainly in two characteristic periods of the freezing process: first, at the very beginning of the initial cooling stage (before 0.1 h), the relative error reaches 1.65%, and then, at the end of the freezing stage (from 4 to 6 h), where the relative errors get to 1.35%. Finally, the relative errors are directly correlated with the freezing rate, which is expected given that the abrupt changes in temperature require an improved temporal discretization. Nevertheless, the relative errors can also be related to the chaotic initial cooling stage of the cabinet which causes slight discrepancies, as illustrated in Figure 12, and with the end of the phase-change process that accumulates the local error previously discussed.
(a) Local evolution of temperatures and their relative errors.
(b) Local evolution of freezing rates and their relative errors.
Figure 13: Comparison of local quantities at three fixed points inside the food between the ground-truth and the reconstructions using thermocamera measurements.
@@ 5.2 Food temperature reconstruction with extrapolated data from the airflow domain
Figure 14: Left: Snapshots of the truncated field of measurements, and of zoomed regions of the ground-truth and the reconstruction fields. Right: Local temperature evolution and its relative error at two fixed points of interest within the food domain.
To test the reconstruction quality of derived quantities in the food, Figure 15 compares the contours of liquid fraction $F_{pc}$ and of volumetric heat capacity $\rho_{s}C_{s}$ at three snapshots between the ground-truth and the inverse framework without measurements in the food domain. The first obvious observation is that the distribution of both derived quantities follows the same pattern described by the isotherms: at the first instant of time, the freezing fronts start to advance from the lateral and upper surfaces, but the latent heat contours do not yet cover large regions of the food; then, at $t=1.5$ h, the freezing fronts concentrate in the center with a slight inclination towards the right lateral surface, while the specific heat contours show clear peaks due to the latent heat; finally, at $t=4$ h, there remains a small area that is not yet frozen, which still concentrates a latent heat peak.
Regarding the reconstruction quality, the results generally indicate that the contours advance at a similar speed but with a slight phase shift throughout the process. In addition, a flatter distortion in the shape of the contours is observed, especially at $t=1.5$ h. Considering the long duration of the process ($t\approx 10$ h) and the nonlinear nature of the temperature and derived quantities evolution, the reconstruction procedure achieves excellent physical representation concerning the ground-truth from the direct model.
Figure 15: Comparison of contours of liquid fraction and volumetric heat capacity at at three snapshots between the ground-truth and the inverse model with extrapolated sensors from the airflow domain.

### v2

@@ 4.3 Greedy algorithm for optimal sensor placement
$\frac{\|T-T^{*}\|}{T}\leq\left(\inf_{c\in\mathbb{R}^{n}}\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{-1}\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}},$ (12)
where $\boldsymbol{A}_{n}$ is the $n$-th low rank approximation of $\boldsymbol{A}$ and ${\|\cdot\|_{\text{FRO}}}$ denotes the Frobenius norm:
@@
The bound (12) shows where the condition $m>n$ comes from, indicating that a reduced model that is too enriched would lead to non-observable modes, a critical issue for well-posedness in data assimilation problems. This provides a criterion for selecting the modes to be used in the inverse method, guarantying that any other configuration for the matrix $\boldsymbol{\Phi}$, would lead to sub-optimal reconstruction results (this does not mean that $\boldsymbol{\Phi}$ always approximates better the forward dynamics as $n$ increases). Using standard arguments, the second term in the right-hand-side of (12) can be expressed in terms of the singular values of $\boldsymbol{A}$:
$\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}}=\frac{\sqrt{\sum_{i=n+1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}{\sqrt{\sum_{i=1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}.$
@@
$\left(\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{2}$ $=\frac{d^{T}\hat{V}^{T}\boldsymbol{G}^{T}\boldsymbol{G}\hat{V}d}{d^{T}V^{T}\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}Vd}=\frac{d^{T}\hat{S}^{2}d}{d^{T}d}=\frac{\sum_{i=1}^{n}d_{i}d_{j}\hat{S}_{i}^{2}}{\sum_{i=1}^{n}d_{i}^{2}},$ (13)
@@
Then, for a given $n$, we aim at selecting the location of $m\geq n$ sensors from a pool of $M$ possible choices over the observable region $\Omega$. The proposed approach follows the underlying idea of a Greedy method, adding the sensors progressively and looking at each stage for the location of a new sensor that yields the smallest a-priori bound (14). Let us recall that, as seen in 4, a measurement sensor can be seen as a linear operation on the temperature field. Thus, considering the Riesz representation of this linear functional, in what follows, we will refer to a sensor as a column of the matrix $\boldsymbol{W}$. Moreover, we will denote as $\mathcal{S}=\{\ell_{1},\ldots,\ell_{M}\}$ a set of all possible sensor locations, and $\hat{\boldsymbol{W}}$ the resulting matrix obtained with the greedy procedure. The step-by-step process is shown in the blue box (Algorithm 1). Besides optimizing the conditioning of the matrix in the normal equations, this approach for the selection of additional modes has the physical meaning of adding experimental measurements at locations where the most relevant dynamics can be observed.
To summarize the overall cost of our procedure, we have gathered the dominant computational costs for every step within the pipeline, as shown in Table 3. The heavy computational burden of the first three steps is kept at the training (or offline) stage, so that whenever a reconstruction must be carried out, only a small number of computations are left, scaling with the number of modes of the SVD. All the steps have behaved well for large scale problems, with the exception of the SVD construction, which could lead to excessive RAM memory consumption. In such cases, the classical method is replaced by the randomized SVD procedure [90, 45], where a stochastic sampling of the column space for the snapshot matrix is carried out, leading to fastest computations, in case it is needed.
Remark: The greedy approach inherits the mathematical analysis of the full method. Let us denote
$\beta(m)=\inf_{c\in\mathbb{R}^{n}}\left(\frac{\boldsymbol{G}c}{\boldsymbol{\Phi}c}\right),$
which controls the stability of the recovery procedure for a given number $m$ of measurements. Observe that our greedy algorithm, by construction (see step 2 of Algorithm 1), is designed so that, if we pick $m^{*}$ measurements with our method, then
$\beta(m^{*})\geq\beta(m_{\text{sub}}),$
for any other selection of sub-optimal measurements $m_{\text{sub}}$, thus leading to the best possible and robust reconstruction for a given number of available measurements. In other words, selecting the sensor that gives the largest possible $\beta(m)$ implies selecting the sensor that gives the smallest possible error according to bound (14).
Table 3: Computational cost of the full pipeline. The free constant $\alpha$ could be up to 2 depending on the iterative scheme used to invert the system of equations and the preconditioning quality. Only the reconstruction and assembly step is done whenever new measurements are received. Taking into account that $n$ is significantly smaller than the full order model degrees of freedom $\mathcal{N}$, we end up with near real-time reconstructions.
Solution
sampling
via FVM
Training
SVD
Computation of
observation
matrix W
Reconstruction with
$n\ll\mathcal{N}$ modes
Assembly
of $u^{*}$
Dominant
computational
cost
$\mathcal{O}(k\mathcal{N}^{\alpha})$ $\mathcal{O}(k^{2}\mathcal{N})$ $\mathcal{O}(m\mathcal{N}^{\alpha})$ $\mathcal{O}(n^{\alpha})$ $\mathcal{O}(n\mathcal{N})$
@@ 5 Numerical examples
The pipeline of the reduced-order inverse estimation framework proposed in this work is illustrated in Fig. 8. In the training stage, we construct a five-dimensional parameter set by gathering snapshots from 64 simulations of the direct model. The resulting time-dependent dataset is obtained by randomly varying the boundary conditions, i.e., $U_{in}$, $T_{C}$, $T_{ext}$, and $h_{ext}$, within a fixed range by using a uniform random distribution, as presented in the lower-left corner of Fig. 8. The magnitude of these variation ranges is consistent with real operating conditions of convective freezers and was estimated from the literature review and technical information from previous works [65, 23]. Then, we train the reduced order space by using an SVD with 48 simulations from the previous dataset, resulting in a total of 1.2×$\times\dot{1}0^{4}$ snapshots. The first 10 modes of the resulting bases are depicted in Fig. 9.a. Thereafter, we manufacture artificial low-quality measurements of the temperature fields from the direct model by degrading the resolution in sensors or pixels of size $2\times 2$ cm2 (see also Fig. 8). Finally, using a Greedy algorithm, we find optimal sensor placements based on a-priori error optimization. In the reconstruction stage, we recover the time-dependent temperature fields from 16 synthetic measurements using direct simulations different from those within the training dataset, using the ROM-based least-squares approach described in the previous section.
For the computational implementation of the numerical experiments, for the finite volume scheme solution of the forward problem, we have used an in-house library called MVF2D [91]. For the full ROM and inverse pipeline, we have used the software MAD ([92, 93], chapter 5), used successfully for forward and inverse problems in previous works such as [56, 94]. The hardware used in the computations for this section were done in a middle-range server using a 48 cores AMD Epyc processor, and 128GB of RAM memory.
Figure 8: Flow diagram of the proposed inverse framework used to reconstruct the temperatures. A training stage is performed to both, sample the solution space with several numerical simulations of the system dynamics (4), and to extract relevant flow features with the ROM basis $\boldsymbol{\Phi}$. Next, temperature measurements are received and assimilated to find their best fit with respect to the column space of $\boldsymbol{\Phi}$
@@ 5.1 Reconstruction of temperature fields from thermocamera measurements
The first test considers full-domain synthetic measurements mimicking a thermocamera placed over the two-dimensional mid-plane of the freezer. This setup results in an image composed of $m=361$ pixels or sensors. Having this setup fixed for the measurement vectors $\ell_{i}$, one can evaluate the matrix conditioning of the normal equations (10) summarized in (14). The decay of the eigenvalues for the covariance matrix associated with the snapshot matrix is shown in Fig. 9. The values and decay tendency of the eigenvalues are related to the ROM’s ability to approximate the original model’s physics. The rapid decay of the values suggests, in particular, that the SVD provides a good approximation of the temperature field in the phase change phenomenon, also with a limited number of modes.
The indicator $e(n)$ from (14) for the fixed measurement setup is shown in Fig. 9$(a)$. Since this quantity is related to an upper bound for the average error when computing the reconstruction, we infer an optimal dimension for the reduced-order model of $n=111$. The curve also clearly shows how adding more modes affects the matrix conditioning, lowering the quality of the expected result due to the reduced observability of the ROM when $n$ approaches $m$.
Figure 9: (a) First 10 SVD basis functions. (b) Eigenvalue decay and a-priori error as the number of the ROM modes $n$ increases. (c) Evolution of the relative $\ell^{2}$ norm error for the 16 simulations out of the training set. We observe an agreement between the theoretical a-priori bound from (14) and the resulting empirical average error for the full test data-set, reaching a 5 % peak at the freezing start and staying below the 1 % until the process ending.
A comparison of the temperature fields at three different snapshots is provided in Fig. 10, showing the ground truth $T_{\text{GT}}$ (not included in the reduced-order model training set), the measurements $\ell=\boldsymbol{W}^{T}T_{\text{GT}}$, and the reconstructions $T^{*}$. As a benchmark criteria, we introduce the relative $\ell^{2}$ error:
@@
which we evaluate for every test case. The average result for the 16 simulations is depicted with respect to time in Fig. 9$(b)$, where we can conclude an overall good reconstruction quality during the whole freezing process, thus validating our framework at a proof-of-concept stage. The reconstruction process and quality depends on how similar the unknown state is to any combination of those in the snapshots dataset. In our case, we train a robust reduced-order model varying the boundary conditions of the forward problem, so that we can ensure a good sensitivity to this parameter.
In addition, we do test exhaustively the reconstruction algorithm using a single, randomly chosen, parameter configuration, with the following values: $U_{in}=0.2\ ms^{-1}$, $T_{C}=-21.7\ \degree C$, $T_{ext}=21.3\ \degree C$, and $h_{ext}=0.88\ Wm^{-2}K^{-1}$. The results are depicted in Fig. 10$(a)$ for 3 relevant time-steps of the full field, and in Fig. 10$(b)$ for a comparison of the liquid fraction contours. We observe a good qualitative agreement between the ground truth and the reconstructed field during the whole freezing process, especially within the food domain, which is of utmost interest for the targeted application. The field reconstruction indicates that the turbulent airflow recirculation through the cabinet is properly described. Moreover, the congruence in the concentration of freezing fronts within the food demonstrates that the phase change process is adequately represented. However, slight differences are noticeable in the upper region of the cabinet, where the airflow is more chaotic due to turbulent fluctuations and the stratification by buoyant forces.
Figure 10: (a) Temperature field evolution of the ground-truth from the direct model, the full-domain measurements, and the reconstruction procedure at three characteristic snapshots. (b) Comparison of contours of liquid fraction and volumetric heat capacity between the forward and inverse results.
@@
$\Phi_{rel,error}(\%)=\frac{|\Phi_{P}-\Phi^{*}_{P}|}{|\overline{\Phi}_{P}|}\times 100.$ (16)
@@
By analyzing the time-dependent relative errors of temperature and freezing rate, we find that the control point in the freezing center presents larger errors than in the food boundaries for the whole process. We must consider that the freezing center represents the point where the freezing fronts concentrate after passing through the entire food domain, so it accumulates all the numerical errors in space and time of the nonlinear food phase-change phenomenon. Therefore, it can be considered that this area is prone to greater numerical errors, which could explain the lower quality reconstruction. Moreover, the local errors increase mainly in two characteristic periods of the freezing process: first, at the very beginning of the initial cooling stage (before 0.1 h), the relative error reaches 1.65%, and then, at the end of the freezing stage (from 4 to 6 h), where the relative errors get to 1.35%. Finally, the relative errors are directly correlated with the freezing rate, which is expected given that the abrupt changes in temperature require an improved temporal discretization. Nevertheless, the relative errors can also be related to the chaotic initial cooling stage of the cabinet which causes slight discrepancies, as illustrated in Fig. 10, and with the end of the phase change process that accumulates the local error previously discussed.
Figure 11: Reconstruction comparison over time for both the temperature (a) and freezing rates (b) at 3 control points. In any of the points, the reconstruction relative error is kept below 2 % for the temperature.
5.2 Reconstruction of the validation test with experimental data
This section discusses a test case to validate the reconstruction capabilities against the experimental data-set of section 3.2, so that we verify that the optimization problem is capable of capturing, with a minimal configuration, the temperature evolution within the control points on the food, where real data is available. To do so, we keep the ROM configuration at its minimum, using the forward simulation from section 3.2 to build up the reduced basis $\boldsymbol{\Phi}$, therefore using a single parameter for this task: time. As with the previous example, we consider square pixels with local averages of the temperature. Following the a-priori analysis made for the previous example, we select the optimal ROM dimension to $n=86$, looking at the minimal value of (14).
[…]
Figure 12: Validation of the inverse framework using test-case with experimental available measurements. (a) The method is able to discover the full temperature field within the working domain from low resolution data, and (b) to fit the empirical data at two control points.
@@ 5.3 Food temperature reconstruction with extrapolated data from the airflow domain
As expected, the quality of the overall reconstruction decreases as the number of sensors is reduced. Despite this, the reconstruction of the temperature fields still achieves great fidelity with respect to the ground-truth; Only slight spatial differences are observed in the freezing fronts within the food. Similar to the previous section, the quantitative differences in temperature evolution between the ground-truth and the reconstruction are hardly distinguishable, except for a short lapse at the end of the freezing stage. Also, the evolution of relative errors shows a similar tendency with two main periods where the errors peak: at the beginning of the initial cooling stage, and at the end of the freezing stage. Nevertheless, when comparing the magnitude of the relative errors with those of Fig. 11, is noticeable that the average errors almost double throughout the entire freezing process, reaching maximum values of 4.8% at the beginning and 4.2% at the end of the freezing stage. The contours of the liquid fraction follow the same pattern described by the isotherms: at the first instant of time, the freezing fronts start to advance from the lateral and upper surfaces, but the latent heat contours do not yet cover large regions of the food; then, at $t=1.5$ h, the freezing fronts concentrate in the center with a slight inclination towards the right lateral surface, while the specific heat contours show clear peaks due to the latent heat; finally, at $t=4$ h, there remains a small area that is not yet frozen, which still concentrates a latent heat peak. Regarding the reconstruction quality, the results generally indicate that the contours advance at a similar speed but with a slight phase shift throughout the process. In addition, a flatter distortion in the shape of the contours is observed, especially at $t=1.5$ h. Considering the long duration of the process ($t\approx 10$ h) and the nonlinear nature of the temperature and derived quantities evolution, the reconstruction procedure achieves excellent physical representation concerning the ground-truth from the forward model.
Figure 13: Depiction of temperature recovery using measurements which are blind to the food region, thus allowing us to extrapolate the field beyond the observed region.(a) Fields comparison in the full domain and in the region near the food at two snapshots. (b) Local temperature evolution and its relative error at two fixed points of interest. (c) Comparison of liquid fraction contours between the forward and inverse models at three snapshots.